# Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria

### v2

@@ Abstract
Large Language Models (LLMs) rely on generating extensive intermediate reasoning units (e.g., tokens, sentences) to enhance final answer quality across a wide range of complex tasks. While generating multiple reasoning paths or iteratively refining rationales proves effective for improving performance, these approaches inevitably result in significantly higher inference costs. In this work, we propose a novel sentence-level rationale reduction training framework that leverages likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences. Unlike previous approaches that utilize token-level reduction, our sentence-level reduction framework maintains model performance while reducing generation length. This preserves the original reasoning abilities of LLMs and achieves an average 17.15% reduction in generation costs across various models and tasks.
@@
Effective Reduction of Redundant Rationale via Principled Criteria
Joonwon Jang1  Jaehee Kim 2  Wonbin Kweon3  Hwanjo Yu1,4,∗ POSTECH1,3,4, Seoul National University2 {kaoara, skwb4453, hwanjoyu}@postech.ac.kr 1,3,4 jaehee_kim@snu.ac.kr 2
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities comparable to human cognitive abilities (Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). These models can solve complex reasoning tasks by explicitly generating extended reasoning paths. The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which additionally enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023). This iterative generation of explicit reasoning paths allows the model to refine and expand its thought processes, incorporating strategic planning and ongoing cognitive generation (Xi et al., 2023; Yang et al., 2024).
@@
Although the extensive generation of explicit reasoning units leads to better peformance, it inherently leads to higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024). Moreover, fine-tuning LLMs using a complete reasoning path does not consistently guarantee stronger performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), suggesting the need for methods that can maintain reasoning performance while reducing the generation of reasoning units. Despite this apparent requirement, it remains underexplored how to maintain LLM reasoning capabilities while reducing intermediate reasoning paths across diverse tasks.
@@
Previous methods (Yu et al., 2024b) have drawn inspiration from human cognitive processes to address the aforementioned limitations. Some studies have proposed training pipelines that utilize augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024). However, they are inherently vulnerable as they significantly depend on the generative capabilities of LLMs. In response, other works have directly trained LLMs without augmenting datasets that facilitate implicit reasoning before answer generation. Deng et al. (2023) introduced a knowledge distillation method to distill explicit reasoning into implicit reasoning through token-level hidden states. Deng et al. (2024b) selected tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens. They progressively eliminate tokens from the beginning of the full rationale during the Chain of Thought (CoT) fine-tuning. However, these methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process. Furthermore, their evaluation has been predominantly conducted on synthetic arithmetic reasoning tasks, which presents limitations for real-world applications.
@@
To address these limitations, we propose a novel training method for maintaining LLM reasoning performance while systematically eliminating redundant reasoning units within the reasoning process. Our method leverages sentences as fundamental reduction units, establishing more linguistically coherent boundaries compared to token-level approaches. We first empirically demonstrate that rationale sentences in early reasoning stages can introduce redundancy in LLM’s answer generation process. Drawing inspiration from Dong et al. (2023), we introduce the concept of ‘verbosity’ and develop a CoT fine-tuning method that removes unnecessary rationale sentences, thereby reducing rationale generation while preserving reasoning capabilities.
The primary contributions of this work are as follows:
  1. 1.
    We propose a novel method for sentence-level reasoning path reduction in LLMs.
  2. 2.
    We demonstrate that principled criteria can effectively identify redundancies and reduce intermediate token generation efficiently.
  3. 3.
    We validate our method’s effectiveness and robustness through extensive experiments across various LLMs and ablation studies.
@@ 2.1 Performance-Cost Tradeoffs in Reasoning Path Generation
Recent research has demonstrated the importance of generating iterative and refined reasoning paths to improve the reasoning capabilities of models, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). Self-Consistency (Wang et al., 2023) improves reasoning accuracy by generating more than 20 distinct reasoning paths and selecting the most consistent answer using ensemble methods, resulting in substantial computational overhead. Similarly, Tree of Thoughts (ToT) (Yao et al., 2023) expands the exploration of reasoning paths using a tree-structured method, while Strategic Chain of Thought (SCoT) (Wang et al., 2024) utilizes an Inner Dialogue Agent to adaptively explore the diverse reasoning path.
Concurrently, Self-Refine (Madaan et al., 2024) utilizes an iterative framework that incorporates feedback to refine responses and enhance reasoning accuracy, necessitating multiple forward passes through the model. Additionally, the Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) guides models to iteratively generate relevant reasoning paths through reflection for accurate answer generation. Given the necessity of iterative generation and refinement of reasoning paths to achieve optimal performance, they inherently increase inference costs and latency. Therefore, it is essential to investigate methods for efficiently generating these paths.
@@ 2.2 Reasoning Path Reduction
To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to shorten the generation of reasoning paths. Yu et al. (2024b) guide model to generate multiple reasoning paths using Self-Consistency to construct datasets, subsequently fine-tuning the model to generate direct answers without intermediate reasoning steps. Liu et al. (2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are subsequently incorporated into the progressive training phase. While these approaches demonstrate empirical effectiveness, they present two fundamental limitations: (1) their heavy dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning path necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
To address these weaknesses, another line of work (Deng et al., 2023, 2024b) has focused on directly training LLM without augmented datasets. Implicit-CoT (Deng et al., 2023) utilizes a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model uses these predicted states to generate answers. ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively removing them from the beginning of the reasoning path within the CoT fine-tuning process. However, these methods have shown limited generalization across diverse datasets as they have only been validated on simple arithmetic reasoning tasks, such as multiplication problems. This limitation raises concerns about their applicability to real-world scenarios where rationales are expressed in natural language. Moreover, they do not justify their criteria for token-level removal nor consider whether more linguistically natural units (e.g., sentences) might serve as more effective reduction units. Specifically, the token-level removal approach may inadvertently eliminate critical information necessary for answer generation or distort the semantic information of the sentence, while requiring repeated removal of redundant reasoning steps. Motivated by the computational costs of the extensive reasoning path generation and the limitations of existing reduction approaches, we investigate the redundancy of various sentence positions for potential removal and propose a novel method with principled criteria that can efficiently reduce reasoning paths while maintaining their effectiveness.
3 Early Stage Rationales are Redundant
Figure 1: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
@@ 3.1 Quanitifying the redundancy
Before delving into the method, we first investigate which positions within the rationale sentences should be selected for reduction. When the likelihood of the answer remains unchanged after a sentence is removed from the full rationales, it indicates that the sentence may be redundant for the reasoning process. To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after the sentence reduction as follows:
@@
3.2 Redundancy of leading sentences
We conduct a pilot study to empirically demonstrate the redundancy of the leading sentences within the rationales. By varying the size of $\{r_{i}\}$ from 1 to 4, we calculate the NLL of Mistral 7B (Jiang et al., 2023) across different reasoning datasets. To explore more general patterns of sentence removal, we employ a stochastic strategy in our experimental setup. We compare three configurations for sampling $\{r_{i}\}$: front, random, and back. In the front configuration, initial sentences are assigned a higher probability when forming $R^{\prime}$, whereas random and back configurations use uniform and progressively increasing probabilities, respectively [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position.. Additionally, we computed the NLL for full rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction. As depicted in Figure 1, the NLL difference between the full and reduced rationales shows a marginal likelihood difference in the front ($<$) configuration. In contrast, removing sentences randomly ($\diamond$) or from the back ($>$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of removal position in the reasoning and answer prediction process (for additional data, see Appendix A).
@@ 4 Verbosity-Aware Rationale Reduction
Figure 2: Overview of our VARR/VARR+ framework. Initially, we select a candidate sentence from the beginning of the rationale (Section 3). After removing the candidate sentence, we evaluate $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$ using Equations (5) and (9). If the candidate sentence meets the verbosity evaluation criteria, it is excluded in subsequent training steps. The model then proceeds with training, where the redundant sentence is excluded from the rationale.
Based on the above observation, we propose the Verbosity-Aware Rationale Reduction (VARR) framework. In Section 4.1, we introduce the concept of ‘verbosity’ as a principled criterion for identifying redundant reasoning sentences. Section 4.2 describes the integration of verbosity in the reduction process during CoT training. Subsequently, Section 4.3 presents additional verbosity term by incorporating the wrong answer to improve robustness. Finally, Section 4.4 illustrates the overall VARR framework.
4.1 Verbosity
To quantify the redundancy of the sentences for removal, we first introduce a foundational concept ‘verbosity’. Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we measure the verbosity of a sentence $r_{i}$ on $y$ by measuring the difference in Kullback–Leibler divergence (KL-divergence) as follows:
@@ 4.1 Verbosity
where $q(y|x)$ is the ground truth distribution. The verbosity$(y)$ highlights the informational contribution or redundancy of a rationale sentence $r_{i}$ on answer $y$. Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), the verbosity$(y)$ can be computed as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 15.00002pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log P(y|R,x)]\\ &\hskip 15.00002pt+E_{q}[\log P(y|R^{\prime},x)]\\ &=\log\left(\frac{P(y_{g}|R^{\prime},x)}{P(y_{g}|R,x)}\right),\end{split}$ (3)
where $y_{g}$ is the ground truth answer (i.e., $q(y_{g}|x)=1$). $H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively. Intuitvely, a higher value of $\textit{verbosity}(y_{g})$ implies that the likelihood of the model generating the ground truth answer increases after removing $r_{i}$, indicating that its removal is beneficial.
@@ 4.2 Verbosity Identification in CoT Training
Given an input sequence, the CoT training (Nye et al., 2021) aims to train LLMs to generate full rationale, followed by the ground truth answer with the following loss:
@@
In each training step $t$, each sentence $r_{i}$ within $R$ is assessed with the following the equation:
@@
Here, we sequentially select $r_{i}$ starting from the first sentence and construct $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, based on our analysis in Section 3, which indicates that early-stage rationales are more likely to be redundant. A non-positive verbosity score indicates that removing $r_{i}$ from $R$ impairs the model’s performance, thereby identifying the sentence as essential and retaining it in the rationale.

### v3

# Verbosity-Aware Rationale Reduction: Sentence-Level Rationale Reduction for Efficient and Effective Reasoning

@@ Abstract
Large Language Models (LLMs) rely on generating extensive intermediate reasoning units (e.g., tokens, sentences) to enhance final answer quality across a wide range of complex tasks. While this approach has proven effective, it inevitably increases substantial inference costs. Previous methods adopting token-level reduction without clear criteria result in poor performance compared to models trained with complete rationale. To address this challenge, we propose a novel sentence-level rationale reduction framework leveraging likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences. Unlike previous approaches, our method leverages verbosity to selectively remove redundant reasoning sentences while preserving reasoning capabilities. Our experimental results across various reasoning tasks demonstrate that our method improves performance by an average of 7.71% while reducing token generation by 19.87% compared to model trained with complete reasoning paths.
@@
Sentence-Level Rationale Reduction for Efficient and Effective Reasoning
Joonwon Jang1,5  Jaehee Kim2  Wonbin Kweon3  Seonghyeon Lee4  Hwanjo Yu1,∗ POSTECH1, Seoul National University2, University of Illinois Urbana-Champaign3 Kyungpook National University4, LG AI Research5 {kaoara, hwanjoyu}@postech.ac.kr jaehee_kim@snu.ac.kr wonbin@illinois.edu sh0416@knu.ac.kr
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities comparable to human cognitive abilities (Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). These works demonstrate the capability to solve complex reasoning tasks through explicitly generating extended reasoning paths. The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which further enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023). Through this iterative generation of explicit reasoning paths, the model refines and expands its thought processes while incorporating strategic planning and continuous cognitive generation (Xi et al., 2023; Yang et al., 2024).
@@
While the extensive generation of explicit reasoning units leads to improved performance, it inevitably results in higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024). Furthermore, fine-tuning LLMs with complete reasoning paths does not necessarily guarantee enhanced performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), indicating the necessity for methods that maintain reasoning capabilities while reducing the generation of reasoning units. Despite this apparent requirement, it remains underexplored how to maintain LLM reasoning capabilities while reducing intermediate reasoning paths across diverse tasks.
Previous methods primarily focused on reducing reasoning paths from two distinct perspectives. Some studies have proposed training pipelines that leverage augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024). However, these approaches remain inherently vulnerable due to their significant dependence on the generative capabilities of LLMs.
@@
In response, other works have focused on directly training LLMs without dataset augmentation to reduce explicit reasoning paths. Deng et al. (2023) introduced a knowledge distillation method to distill explicit reasoning into implicit reasoning through token-level hidden states. Deng et al. (2024b) adopted tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens while Hao et al. (2024) compressed complete rationales into the predefined number of hidden states of tokens. However, their methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process. Moreover, their evaluation has primarily focused on synthetic arithmetic reasoning tasks, limiting their applicability to real-world scenarios.
@@
To address these limitations, we propose a novel training method that maintains LLM reasoning performance while systematically reducing redundant reasoning units within the reasoning process. Our method adopts sentences as fundamental reduction units, establishing more linguistically meaningful boundaries compared to token-level approaches. Through empirical analysis, we demonstrate that sentences in early rationale steps can introduce redundancy in the LLM’s answer generation process. Inspired by Dong et al. (2023), we introduce the concept of ‘verbosity’, a likelihood-based criteria, to identify redundant reasoning sentences. By incorporating verbosity identification into the training process, the model excludes redundant reasoning sentences, thereby reducing intermediate token generation. Finally, we demonstrate our method’s effectiveness and generalizability across various real-world reasoning datasets, showing our method improves performance by an average of 7.71% while reducing token generation by 19.87% across various LLMs, and through systematic ablation studies, we analyze the contribution of each proposed component.
@@ 2.1 Performance-Cost Tradeoffs in Reasoning Path Generation
Recent research has demonstrated the critical role of generating iterative and refined reasoning paths in enhancing model reasoning capabilities, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). Self-Consistency (Wang et al., 2023), Tree of Thoughts (ToT) (Yao et al., 2023), and Strategic Chain of Thought (SCoT) (Wang et al., 2024) improve reasoning accuracy through ensemble-based path selection, tree-structured exploration, and adaptive reasoning with an Inner Dialogue Agent, where each approach requires iterative reasoning path generation, resulting in substantial computational overhead.
Concurrently Self-Refine (Madaan et al., 2024) and Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) enhance reasoning abilities through iterative feedback-based refinement and reflective path generation, respectively, though both require multiple forward passes through the model. While the iterative generation and refinement of reasoning paths are essential for achieving optimal performance, they inherently increase inference costs and latency. Therefore, it is crucial to investigate methods for efficiently generating these paths.
@@ 2.2 Reasoning Path Reduction
To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to reduce the generation of reasoning paths. Yu et al. (2024b) employs Self-Consistency to generate multiple reasoning paths for dataset augmentation, then fine-tunes the model to produce direct answers. Liu et al. (2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are then integrated into the progressive training phase. While these approaches demonstrate empirical effectiveness, they exhibit two fundamental limitations: (1) their substantial dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning paths necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
To address these weaknesses, another line of work (Deng et al., 2023, 2024b; Hao et al., 2024) has focused on directly training LLMs without augmented datasets. Implicit-CoT (Deng et al., 2023) implements a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model leverages these predicted states to generate answers. ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively eliminating them from the beginning of the reasoning path within the CoT fine-tuning process. However, these methods demonstrate limited generalization across diverse datasets as they have been validated exclusively on simple arithmetic reasoning tasks, such as multiplication problems. This limitation raises concerns about their applicability to real-world scenarios where rationales are expressed in natural language. Furthermore, they do not explore the adoption of principled criteria and linguistically natural units (e.g., sentences). Specifically, the token-level reduction approach may eliminate critical information necessary for answer generation or distort the semantic information of the sentence. Motivated by these limitations in existing reduction approaches, we examine the redundancy of various sentence positions for potential elimination and propose a novel method with principled criteria that can effectively reduce reasoning paths while maintaining their efficacy.
3 Early Step Rationales are Redundant
Figure 1: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL with the complete rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
@@ 3.1 Quanitifying the Redundancy
Before delving into the method, we first investigate which positions within the rationale sentences should be selected for reduction. When the likelihood of the answer remains unchanged after removing a sentence from the full rationales, this indicates that the sentence may be redundant in the reasoning process. To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after sentence reduction as follows:
@@
3.2 Redundancy of Early Reasoning Sentences
We performed a pilot study to empirically demonstrate the redundancy of leading sentences within the rationales by analyzing the NLL of Mistral 7B (Jiang et al., 2023) across diverse reasoning datasets. Specifically, we varied the size of $\{r_{i}\}$ from 1 to 4 and investigated general patterns of sentence removal using a stochastic approach. To compare different sentence selection configurations, we considered three sampling methods for $\{r_{i}\}$: front, where initial sentences were prioritized; random, with uniform probabilities; and back, where probabilities progressively increased for later sentences [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position.. Additionally, we computed the NLL for complete rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction. As illustrated in Figure 1, the front ($\triangleleft$) configuration shows only marginal NLL differences relative to complete rationale sentences. In contrast, removing sentences randomly ($\diamond$) or from the back ($\triangleright$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of a candidate rationale position strategy for removal in the reasoning and answer prediction process (for additional analysis, see Appendix A).
@@ 4 Verbosity-Aware Rationale Reduction
Figure 2: Overview of our VARR/VARR+ framework. Initially, we select a candidate sentence from the beginning of the rationale (Section 3). After selecting the candidate sentence, we evaluate Equations (5) and (9) by calculating $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$. If the candidate sentence meets the verbosity evaluation criteria, it is excluded in subsequent training steps. The model then proceeds with training, where the redundant sentence is excluded from the rationale.
Based on these observations, we propose the Verbosity-Aware Rationale Reduction (VARR) framework. In Section 4.1, we introduce the concept of ‘verbosity’ as a principled criterion for identifying redundant reasoning sentences. Section 4.2 elaborates on how we integrate verbosity into the reduction process during CoT training. In Section 4.3, we extend the verbosity term by incorporating incorrect answers to enhance robustness. Finally, Section 4.4 presents the comprehensive VARR framework.
4.1 Verbosity as Principled Criterion
To quantify the redundancy of sentences for potential removal, we introduce the fundamental concept ‘verbosity’. Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we quantify the verbosity of a sentence $r_{i}$ on $y$ by computing the difference in Kullback–Leibler divergence (KL-divergence) as follows:
@@ 4.1 Verbosity as Principled Criterion
where $q(y|x)$ is the ground truth distribution. The verbosity$(y)$ measures the informational contribution or redundancy of a rationale sentence $r_{i}$ with respect to answer $y$. Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), we can express the verbosity$(y)$ as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 6.99997pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log p_{\theta}(y|R,x)]\\ &\hskip 6.99997pt+E_{q}[\log p_{\theta}(y|R^{\prime},x)]\\ &=\log\left(\frac{p_{\theta}(y_{g}|R^{\prime},x)}{p_{\theta}(y_{g}|R,x)}\right),\end{split}$ (3)
where $y_{g}$ denotes the ground truth answer (i.e., $q(y_{g}|x)=1$). $H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively [^2] For the sake of explainability, we assume each expression’s $y$ is represented by a single token.. Intuitvely, a higher value of $\textit{verbosity}(y_{g})$ implies that the likelihood of the model generating the ground truth answer increases after removing $r_{i}$, indicating that its removal is beneficial.
@@ 4.2 Verbosity Identification in CoT Training
Given an input sequence, CoT training (Nye et al., 2021) aims to train LLMs to generate complete rationale, followed by the ground truth answer:
@@
During each training step $t$, we evaluate each sentence $r_{i}$ within $R$ using the following criterion:
@@
Here, we sequentially select $r_{i}$ starting from the first sentence and construct $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, based on our analysis in Section 3, which indicates that early-stage rationales are more likely to be redundant. When the verbosity score is non-positive, it indicates that removing $r_{i}$ from $R$ would impair the model’s performance, thus identifying the sentence as essential and preserving it in the rationale. Detailed training procedures will be described in Section 4.4.